OpenStudy (mathisfun13):

Find the distance from point A (4,2) to point B (-3,2) (answer) units

OpenStudy (anonymous):

The equation that is used for finding the distance (D) between two points is: . . Now let's identify the two points you were given. Let's call point A (4,-3) point 1 and that means that x1 = 4 and y1 = -3. That means that point B (-4, 3) is point 2. Therefore, x2 = -4 and y2 = 3. (If we had let point A be point 2 and point B be point 1, it wouldn't make any difference in the answer we get for the distance. It would still work out the same as below even though the values of x1, y1, x2, and y2 would be changed.) . Now all we have to do is substitute our values for x1, x2, y1, and y2 into the distance equation. When we do we get: . . This simplifies to: . d=sqrt(-8)^2+(6)^2

OpenStudy (mathisfun13):

i dont get it

OpenStudy (mathisfun13):

what does this sign ------> ^ mean?

OpenStudy (anonymous):

^ means power of lets say 2^2=4 :) and 3^2=9

OpenStudy (mathisfun13):

oohh so it would be 64 & 36?

OpenStudy (anonymous):

The equation that is used for finding the distance (D) between two points is: d=sqrt(x2-x1)^2+(y2-y1)^2 u need to identify what is ur x2,x1,y2,y1 after u identify them, then apply the formula d=sqrt(x2-x1)^2+(y2-y1)^2 d=sqrt(-4-(4))^2+(3-(-3))^2 yeah u got 64 and 36 :)

ganeshie8 (ganeshie8):

A (4,2) to point B (-3,2) ^ ^

OpenStudy (mathisfun13):

yea

ganeshie8 (ganeshie8):

distance between points = \(\sqrt{(x2-x1)^2 + (y2-y1)^2}\) \(\sqrt{(4--3)^2 + (2-2)^2}\)

ganeshie8 (ganeshie8):

try to simplify...

OpenStudy (mathisfun13):

so √(7)2 + (0)2

ganeshie8 (ganeshie8):

yes, keep going

OpenStudy (mathisfun13):

so then 49+0=49 units

ganeshie8 (ganeshie8):

square root of that.

ganeshie8 (ganeshie8):

why u always ignoring squareroot ha...

ganeshie8 (ganeshie8):

\(\sqrt{49}\) is not same as \(49\)

OpenStudy (mathisfun13):

lol ohhh 7 units?

ganeshie8 (ganeshie8):

Yes ! good work !

OpenStudy (mathisfun13):

thanksss! your the best on here

ganeshie8 (ganeshie8):

awww ty lol :)